# Comparison of the Migdal transition probabilities in electron-atom inelastic cross sections

### v1

@@ 6 Discussion and conclusion
The excitation/ionization probability from dark matter-nucleus scattering is compared to the electron-atom inelastic cross section. The errors are $\sim 20\ \%$ for Ar at 10 keV but large for Xe compared to the database value, especially excitation shown in Fig. 1. This result may imply that the transition probability of Migdal effect need to be modified for Xe atom. Actually, the calculated bound energies are reported to have $\mathcal{O}(10)\ \%$ error with experimental data Ibe:2017yqa . Note that this result does not compare the transition probability from the inner shell electrons because the probability is dominated by the outermost electrons.
The use of screened Coulomb potential does not show the drastic improvement of their correspondence. But, the additional parameter of screening radius improved the excitation probabilities. For ionization, the simple Coulomb potential shows the better agreement with the data by comparing Fig. 1 and Fig. 2.

### v3

@@ 5 Discussion and conclusion
The excitation/ionization probability from dark matter-nucleus scattering is compared to the electron-atom inelastic cross section. The errors are $\sim 20\ \%$ for Ar at 10 keV and $\sim 60\ \%$ for Xe compared to the database value shown in Fig. 1. For Xe, the database value is limited and the error will be improved for higher energy because the plane wave and the Coulomb wave approximation used for $u$-channel processes becomes better for high energy. Also, the approximation is not good for excitation cross sections because the excited states are approximated to the plane wave states. Therefore, the expected errors for excitation are not reliable. The differences of threshold energy comes from the error of calculated bound energy. Actually, the calculated bound energies are reported to have $\mathcal{O}(10)\ \%$ error with experimental data Ibe:2017yqa . Note that this result does not compare the transition probability from the inner shell electrons because the probability is dominated by the outermost electrons.